The product of 7 and 6¾ is (a) 42¼ (b) 47¼ (c) 42¾ (d) 47¾ 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (b) is correct here
First we have to convert the mixed fraction into improper fraction 63⁄4 = 27/4

= 7 × (27/4)
= 189/4
= 47 whole 1⁄4


Thanks